• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Lindgren

    Log entry time 08:05:19 on March 3,2009

    Entry number 264276

    keyword=Owl Shift Summary

    Owl Shuift Summary

    Owl Shift starting
    Sl Rcihard Lindgren
    TO Vladimir Nelyubin
    3rd Miha Mihavilovic


    00:15 MCC reports that cryolevels 85% filled and soon will try to put beam inti North Linac. Probably take an hour. If all goes well, we will have beam back by 2:00 optimistically.

    3:30 MCC called and reported that they are still tuning cavities trying to establisg a gradient. At least another hour before any beam in machine.

    5:40 MCC reports that NLO4 cryomodule overfilled recylcing
    beam expected by 8:00 AM

    No beam on Owl Shift.